OpenStudy (anonymous):

Prove that for all numbers a and b, \[|a+b| \leq |a| + |b|\]Prove that for all numbers a and b, \[|a+b| \leq |a| + |b|\]@Mathematics

OpenStudy (sasogeek):

dear math, please grow up and solve your own problems! :P

OpenStudy (asnaseer):

just consider each of these cases in turn to get your proof: a<0, b<0 a<0, b=0 a<0, b>0 a=0, b<0 a=0, b=0 a=0, b>0 a>0, b<0 a>0, b=0 a>0, b>0

OpenStudy (anonymous):

the book said you only need 4 cases, and not nine :-D

OpenStudy (asnaseer):

true, you could make use of the fact that a+b=b+a to eliminate some

OpenStudy (asnaseer):

by the way, I was just wondering if there some story behind why you chose the name @agdgdgdgwngo?

OpenStudy (anonymous):

consider the case that a geq 0 and b geq 0, it follows that |a + b| = a + b, which is equal to |a| + |b| consider the case that a leq 0 and b leq 0. it follows that |a + b| = -(a + b) = -a + (-b) = |a| + |b| consider the case that a geq 0 and b leq 0, hence, we must prove that |a + b| leq a - b there are two subcases here: a - b geq 0 or a - b leq 0 for the first subcase we can show that a + b leq a - b by simply subtracting a, proving that -b geq b for the second subcase we can show that -a - b leq a - b, and by simply adding b, proving that a geq -a same thing with the case that a leq 0 and b geq 0, but with a and b switched. tombstone

OpenStudy (asnaseer):

yup - so 4 cases is indeed necessary and sufficient

OpenStudy (anonymous):

:( silly book. the next page says there is a much easier proof

OpenStudy (asnaseer):

I guess the corollary to "think ahead" is "read ahead" in this case! :-)

OpenStudy (anonymous):

alright here is the books new and improved proof (I won't be surprised to see an even easier proof 400 pages later though): using |a| = sqrt{a^2}, (|a + b|)^2 = (a+b)^2 = a^2 + 2ab + b^2 leq a^2 + 2|a|*|b| + b^2 = |a|^2 + 2|a|*|b| + |b|^2 = (|a| + |b|)^2 now root both sides to get that |a+b| leq |a| + |b| since if x^2 < y^2 then x < y provided that x geq 0 and y geq 0 (the book says I have to prove that soon) :(

OpenStudy (asnaseer):

wow - neat proof